Rachelle spends 330 min/wk exercising. Her ratio of time spent on aerobics to time spent on weight training is 6 to 5. How many minutes per week does she spend on aerobics? Howmany miutes per week does she spend on weight training? (could you also explain how you got the answer?)
Answer by stanbon(75887)

Rachelle spends 330 min/wk exercising.
Her ratio of time spent on aerobics to time spent on weight training is 6 to 5.
------
6 to 5 is the same as 6x to 5x, where 6x is aerobic time and 5x is weight
training time.
-------------
How many minutes per week does she spend on aerobics? How many minutes per week does she spend on weight training?
------
Equation:
6x + 5x = 330 min
11x = 330
x = 30 min
----
6x = 6*30 = 180 min (aerobic time)
---
5x = 5*30 = 150 min (exercise time)
====================================
Cheers,
Stan H.
===============